SBS DAQ meeting summaries

From Hall A Wiki
Jump to: navigation, search

SBS DAQ Meeting template


June 1st 2024

May 13th 2024

April 29th 2024

March 25th 2024

March 11th 2024

February 26th 2024

February 19th 2024

February 5th 2024

January 29th 2024

January 22th 2024

December 18th 2023

November 13th 2023

August 21st 2023

August 14th 2023

July 10th 2023

June 12th 2023

June 5th 2023

May 22nd 2023

May 1th 2023

April 11th 2023

April 3rd 2023

February 8th 2023

February 2nd 2023

December 14th 2022

November 9th 2022

October 19th 2022

October 12th 2022

October 5th 2022

September 14th 2022

August 31st 2022

August 24th 2022

August 3rd 2022

July 27th 2022

July 20th 2022

July 6th 2022

June 29th 2022

June 1st 2022

May 25th 2022

May 4th 2022

April 14th 2022

April 1st 2022

March 25th 2022

March 18th 2022

March 11th 2022

March 4th 2022

February 25th 2022

February 4th 2022

January 28st 2022

January 21st 2022

January 14th 2022

January 7th 2022

December 17th 2021

December 10th 2021

December 3rd 2021

November 18th 2021

November 11th 2021

November 5th 2021

Octorber 29nd 2021

Octorber 22nd 2021

October 15th 2021

October 8th 2021

October 1st 2021

September 24th 2021

September 17th 2021

September 10th 2021

September 3rd 2021

August 27th 2021

August 20th 2021

August 13th 2021

August 6th 2021

July 30th 2021

July 23rd 2021

July 16th 2021

July 2nd 2021

June 25th 2021

June 17th 2021

June 11th 2021

June 4th 2021

May 28th 2021

May 21st 2021

May 14th 2021

May 7th 2021

April 30th 2021

April 23rd 2021

April 16th 2021

April 9nd 2021

April 2nd 2021

March 26th 2021

March 19th 2021

March 12th 2021

March 5th 2021

February 26th 2021

February 12th 2021

February 5th 2021

January 29th 2021

January 15th 2021

January 8th 2021

December 18th 2020

December 11th 2020

December 4th 2020

November 20th 2020

November 13th 2020

November 6th 2020

October 23rd 2020

  • Updates
    • HCal
      • work on LED system
      • Bus error still occuring, will try with just FADC
      • Start and run scripts
    • Hodoscope
      • still being repaired 20 bars ready and being stacked
      • Mark starting to
    • Shower
      • one bad phototube, have to unstack
      • setup CODA3 on tedbbdaq
      • Second crate VXS
      • WAPP trigger : shower and preshower - low and high threshold for pions - runs right after GEn RP
  • Setup trigger superviser
    • CDet
      • one pixel pedestal shifting from run to run
  • Ref channel using RF time -
    • MPD with 41.6 MHz, put scintillator in FADC - look at timestamp to correct jitter -
    • plan send RF time to 1 FADC channel
    • need to check software
  • LHRS : trigger HRS for HCal efficiency - LHRS selecting pion for calibration - timing of HRS trigger -
    • upgrade new TI and intel CPU
    • LHRS : need two fiber port - 4 fibers to LHRS trigger - multimode fiber - need new TI transceivers
    • VETO scheme with TI
    • estimate timing

October 16th 2020

  • Updates
    • HCal
      • taking cosmics, warning bus error but data looks ok. CODA3 . Should add
      • need FADC for reference time ( on loan to Steve )

    • Hodoscope : document submitted, hodoscope repairs doing good - aim to take cosmics in November - need to modify code from V1190 -
      • 5 Hall C FADC - Brad can loan 4 FADCs for hodoscope temporarly -
      • can use for beginning
      • Reference time for V1190
    • Calorimeter
      • received quote for FADC
      • VXS crate in TEDF
      • ask for FADC from Chris can get 8 trying to get 10
    • GEM DAQ
      • issues on readout, one MPD in slot 7 will be replaced - to try
      • Paolo check for production of new MPDs, all components for 10 to 12 MPDs available, delay chip on MPD is CERN chip, 7 only but try to get a few more should be ok. Improvement in MPD could be implemented. Need to let Paolo know when MPD production happen he can start working that.
      • multifiber cable  : check cost

    • CDet
      • computer moved to Test Lab for CODA3 developement
      • asking to remove one or 2 fastbus and setup VETROC in VME64X crate
    • LHRS
      • upgrading to CODA3
      • update to intel CPU or CODA3 for vxworks ?
      • Jitter and delay from new TS -
      • Calibration run : trigger ? HRS left trigger ?
      • need to think about fiber going to LHRS

    • Reference time

October 2nd 2020

  • Updates
    • HCAL
      • DAQ back to normal , all FADC and F1 working
      • Added switch
      • Check HV Rpi
      • FADCs using the VXS SD
      • F1 unplugged sync and get trigger from bottom crate
      • latency in second crate : warning latency greater than window, might need to adjust delay and latency
      • both VTPs are booted up : can log in using sbs-aonl account
      • need to add transceivers and fibers
    • GEM
      • UVA GEM : taking cosmics with data no more issues
      • Bigbite GEM will be set up in the TEDF next month : optical readout
      • INFN GEM : Ezekiel restarted HV and DAQ in INFN clean room, HV fine, same issues as before MPD7 has 1 backplane not initializing, now disabled. MPD 17 an 18 have issues ( last ones ) , low level look ok , but with CODA run time out on MPD 18, after disabling APV on MPD 18, time out on MPD 17. Cosmics data taking on going.

Issues with HV and CODA, when kill CODA looses Rpi connection. Might be a java client ? Need to check kcoda scripts. MPD 17 and 18 worked at some point.

      • Paolo working on the firmware with Ben's code
      • GEn RP : barely enough MPDs for spare. Need MPD design ? If less than 10 look for older FPGAs. Money for 4 or 5 for Moller. How many spare MPDs ? HU might have a few spare. 110 APV to make spare.

Inventory of MPDs. Could fix issue of optical link if make a new production ? Check with Dustin and Michael how many spares we could get.

      • GEM VTP readout to be decided
      • INFN clean room : 220 Hz at 100 % occupancy, after MPD buffering x2 or 3 improvement in rate so around 600 Hz at 100 % occupancy.

    • Bigbite
      • Shower : Chuck removed spare Fastbus, TS crate might have been moved, looking for FADC crate for Shower

September 25th 2020

  • Updates
    • HCAL
      • check cabling, Juan Carlos arriving on Sunday
      • F1 no trigger
      • VTP and SD added, serial portserver set up
      • sbsvtp1 has issue booting being , sbsvtp2 SD updated and booted, check boot from atedf3
    • CDet
      • Peter was able to access the test lab and checked set up, detector working as before
      • Start working on VETROC with David
      • VETROC sole source submitted for Connecticut order
      • need take one Fastbus out to install VXS crate
    • GEM
      • UVA clean room ok after TI fix, need to fix APV errors, some MPDs need to be reset from time to time around every 10 k events, might be fixed when all APV errors are cleared, and might be sensitive to noise. Module readout when ready, if readout fails then reset
      • INFN : Ben send code for new firmware to Paolo, Paolo start to simplify code.

Need to start using zero supression code. Xinzhan started looking at zero supression code, will coordinate with Bryan and Hanjie to generate the zero supression parameter. Code should be able to work from either SSP or VME. One file from Bryan available, can run code on this file.

    • Bigbite
      • still taking cosmics, a few channels need to be investigated
      • will set up VXS crate with a few FADC, need BNC to lemo cable, need 243 cables

remove Fastbus and set up VXS crate ( Mark will check with Chuck )

      • VXS TS move ( in rack of Fastbus )

DAQ September 25th 2020

September 18th 2020

  • Updates
    • HCAL
      • Scott and Juan Carlos got CODA3 readout list working , could look at data with SBS offline ( need to be careful that crate ID could change )
      • VXS crate was swapped. Need check F1 trigger ( might be missing ), can read the two FADCs.
      • Asked Ben for a VTP and SD, need a terminal server for VTP serial setup. Should be able to have them setup next week
    • INFN cleanroom
      • 14 MPDs readout with SSP, each has 9 to 15 APV connected. When running in buffering mode, one trigger get lost, investigating if it is coming from one MPD and might need set up busy trigger rules on TI. Right now 200 Hz trigger.

Ben working on MPD simulation to try to determine if that comes from MPD.

      • trigger lost on MPD 7 and last slot with regular VME when Evaristo was testing. Other APV cannot be configured and were taken out from config. Bryan will give more informations on errors to Evaristo for cross check.
      • cosmics data will be taken when Ezekiel will be able to take more data. Maybe check first everything is working with VME, but optical should work, Bryan made a config file working for both VME and SSP
    • UVA GEM cleanroom:
      • Issue with trigger distribution, likely from TI master. Could use regular TI in VXS crate if cannot fix easily. Need one with 3 ports. Will get another TI set up with 3 ports.

    • Bigbite :
    • Shower still taking cosmics , a few miscabling found which need to be fixed
      • VXS crate for TS could be moved to a more convenient place
      • FADC crate ordered, will try to borrow VXS crate and FADC so we can set up
    • Hodoscope:
      • Mark will look into setting the CAEN ADC with a pulser

    • CDet
      • no update but should be able to work when can access next week to get cosmics going again

September 11th 2020

  • Updates
    • HCAL : Bryan looked at crashing issue and sent changes to make to the readout list, will be implemented ( until Wednesday )
    • Hodoscope : Ralph focusing on hodoscope repairs. Document not started, if we have a template document for another detector that could be helpful
    • Bigbite :
      • cosmics data taking - still need to look more into the data -
      • all ADCs cables
    • CDet
      • Ralph was able to run DAQ, cosmics data taken, need to look at data
      • CODA3 setup : waiting for issue with Java and CODA3 - Bryan sent a tar ball with Java ***sbs2 only has 32 bit OS so issue running CODA3 - need to look if need to upgrade OS or find a new computer - will use Duke computer and Hall C
  • Bryan successfuly managed to use the cheap ATOM processor Concurrent controller, so good option if no 2eSST needed and no high performance required ( like VTP readout )
      • VXS crate will be moved from ESB to test lab with Duke computer - can timeshare for VETROC developement
    • INFN GEM optical setup : could read out MPD using VME with firmware and CODA3 - a few APVs not on or working on slot 17 and 18. Need to check cabling

Almost ready for SSP readout : debugging optical system

      • Ben will send common noise zero suppression today or Monday
      • SSP : might need the DDR module

September 4th 2020

  • HCAL
    • CODA3 readout working for VXS crate, but crashes for the VME64X when trying to read the TI information.

Bryan will have a look.

  • APV25
    • Paolo waiting for Ben's code
    • INFN clean room : SSP readout setup, need to do MPD configuration with VME, Ben upgraded MPD to latest firmware
    • UVA cleanroom : issue with TI, need to power cycle the crate and maybe replace the TI
  • CDet
    • turned on breaker, crates turned on, Ralph was able to take data. Seems ready to take cosmics data. Peter might have access to Test Lad during SAD
    • David still working on CODA3, get out of memory when starting java, CODA event builder will crash if less than 16 GB or less. Problem with downloading more than once, might find a workaround, should work with the correct version of java.
  • TDIS meeting DAQ, every month Thursdays 1 PM first week of the months ( except next month ). Cerenkov trigger with VETROC and more FADCs for LAC.

August 28th 2020

  • Updates
    • HCal : still testing readout list for CODA3,

upgrading firmware of TI. Crashed after running library test

      • need setup remote power cycling
    • Bigbite Shower : still taking cosmics data
    • GEM INFN : VXS crate with SSP, issue with CPU will replace it. Ben will update MPD firmware on sbsvme20

In UVA clean room : tested lower speed and setup was stable. Paolo back from vacation, Ben will send SSP code for common mode subtraction so he can try to fit on MPD

Bryan, is installing CODA3 on triton, will finish when other CPU in

    • CDet : Ralph on site at JLab working on hodoscope, can try to get the CDet cosmics going again. Peter will have a video meeting with Ralph to try to restart setup. Could try CODA3 with that cosmics setup. CNU starts in person teaching next week, hope Kara can get back to DAQ.

August 21st 2020

  • Updates
    • HCAL
      • CODA 3 readout list done, need to turn on pulser to get trigger to test
    • CDet
      • Kara coming back to campus next week

VETROC quotes for Catagna and Connecticut ready, will start on purchase request

      • David still working on CODA3 for CDet
  • GEM INFN : will try to setup optical readout with

INFN electronics so Ben can continue testing easily Ben tried lower from 2.5 to 1.25 Gbps. Max trigger 2.7 KHz, Ben think he can fit the common noise supression on MPD which would bring it back to at 5.4 KHz or more depending on occupancy. Will try to set up meeting with Paolo.

  • Bigbite shower : Mark was able to take data at different voltages
  • Hodoscope : Ralph working on write up defining DAQ and software and will work on software later. Ralph approved on site can come next Monday

August 14th 2020

  • Updates
    • HCAL : CODA3 setup , need to move the old read out list to CODA3
    • CDet : CODA3 setup by David, still working on read out list, can meet with Bryan later, gather electronics for CDet using VXS crate. Working on quote and paperwork for VETROC order by Catagna and UConn
    • Bigite : investigated DAQ issues, bad version of java, solved with reverting to java, check with THia to order VXS for shower
    • GEM : Ben discussed trying lower speed - Plan to setup optical link in INFN clean room

July 31st 2020

  • Updates
    • MPD issues with high speed : will discuss it at GEM meeting on Wednesday
    • Bigbite : need to make sure all sessions are correct, will try again
 Glasgow could fund 6 to 8 FADCs - Got quote from Chris Cuevas : 6.2 K$ for 16 units and 5.9K for 25 units about 48 K$
    • CDet : bid process was finished, will start ordering VETROC - Catagna is buying the VXS crate
    • GRINCH : can log in the computer - still need to turn on detector - GRINCH is open right now

July 24th 2020

  • Updates
    • APV25 SSP setup : Ben debugging the system, optical signal looks ok

so maybe SSP firmware bug.

8 to 9 KHz

Jitter on signal might be because of clock power noisy

might be able to modify board

    • HCAL : enpcamsonne, CODA 2.6 still running, can stop them for CODA install, Bryan installed JAVA required, so ready to install CODA3. Bryan updated scripts so they can easily be ported to other hosts.
    • Bigbite : CODA can working, need to look at setup for trigger - Add vncserver to tedbbdaq
    • CDet : David trying to install CODA3

July 10th 2020

  • HCAL : can use DVCS pulser ? Can give delay
        3 outputs from TI to control the LED
        Good time to put VXS crate ? 
        DAQ running 

  • GEM : Xinzhan still check for status, checking the code
      Ben : test in clean room, low voltage turned off
  • Bigbite : still need to check if DAQ running after it was turned on, Bryan has all informations for shower DAQ
  • GRINCH : still cannot connect to VME, will power cycle it
  • Setup remote power cycling
  • HCAL timing : need timing scheme for coincidence HCAL Bigbite
  • Hodoscope : 64 channels V792 for calibration runs

June 26th 2020

  • APV25 GEM update
    • Xinzhan working with Danning, submitted access request last week
  • Bigbite : try to reboot everything so can work remotely

need to send busy from V792

  • CDet : write-up by Kara
  • Plan for GEM with Nilanga : Krishna Kumar can help with GEM DAQ , Chandan could help with design of shielding hut, Hanie could work on BigBite GEM DAQ in TEDF

Need electronics for UV chamber, 10 K$ of APV card ( 150 chips left at UVA)

Most likely INFN + UV not fit in stack.

For UV readout need APV25 or adapter

For next week : Turn/reboot TEDF crates 3r DAQ GEM for GEM UMass postdoc for shield house and BigBite GEM DAQ

June 19th 2020

  • VETROC discussion and introduction. VETROC can take any differential signal so existing.

Existing current firmware only uses front pannel : good enough for testing. Will eventually merge triggered firmware and high resolution firmware. Once ready will switch to it. No student allowed at JLab will try to setup a VME VETROC DAQ system at CNU.

VETROC manual at :

  • BPM/BCM : short discussion
  • HCAL : need to turn on crates so can send readout list to Bryan
  • Hodoscope : same as HCAL

June 12th 2020

First meeting after going back to Medcon5


1) Near term BB DAQ to support the PReshower/Shower after install of new preshower blocks.

2) Setting up the final BB DAQ with integration of all detectors.

  • Shower
    • Need to add sums and TDC informations
    • 1 week of cosmics before finalizing DAQ
    • finalize DAQ after ( CODA3 - TS ) Let Bryan know the computer for CODA3
  • Timing hodoscope
  • Timeline
    • GRINCH : update to CODA3 first
    • Shower Fastbus
    • Hodoscope
    • GEM DAQ : 1 layer soon - need GEM 3rd DAQ setup ? INFN DAQ moved to TEDF

all subsystems give info to Bryan

3) Replacing the CDET fastbus TDc setup with an VME based TDC setup.

4) HCAL plan : switch VME64X to VXS

2 weeks of data - after switch CODA3 and VXS - addditionnal FADC for reference channel

5) Beamline readout for SBS : scalers for BCM, FADC for BPMs / raster , try digital readout in parallel ?

Use event blocking ?

If large event size - event blocking might not help

Readout of scalers - decoding might get broken -


March 27th 2020

First meeting after JLAB to Medcon6

  • Status
    • HCAL

Scott looking at g4sbs simulation. Everything is off, besides enpcamsonne for files. Might install CODA3. Implementing different timing measurement : 1/4 peak height might not work for cosmics ( too small ), might try implement constant fraction algorithm.

    • CDet : HV off , PS Nino off, Not sure if Fastbus is off, computer sbs2 still on for analysis.
    • Bigbite : everything is off besides computer tedbbdaq. Not sure about Grinch. Had to turn on network switch
    • Hodoscope: crate off - maybe work om software analysis - add V792 for calibration - Rachel will coordinate with
    • GEM UVA : All electronics in EEL cleanroom off - computer still on - eel124gemdaq . Not sure if silo working.
  • Task that can be done remotely
    • mostly CODA3 install and decoding analysis software

March 13th 2020

  • Status subsystems:
    • HCAL ( Scott ) : timing studies pdf - maybe cable length
    • switch to VXS crate - PMT greasing - needs to know if Juan Carlos can help
    • HCAL CODA crashing almost every end of run : maybe loss of synchronization - errors message:

Bogdan : possible AC pickup - check with LED

    • CDET: start running test of Fastbus TDC, one crate for cosmics and one for testing. CNU cancels face to face courses for next 4 weeks
    • GEM readout : issues on MPD, fixed firmware, up to 8 units, Danning continue scaling to 32 MPDs

February 14th 2020

  • HCAL : Scott still working on timing resolution, reach 0.5 ns for most tubes, FADC timing resolution at 2.5 ns level, should be able to be improved

February 7th 2020

  • Updates
    • HCAL status( slides by Scott pdf )

Timing resolution about 0.8 ns, time walk correction done, JLab tubes don't have enough gain : change range to 0.5V for FADC, can add amplifier, estimation of the number of photoelectrons ( maybe around 5 to 8 per MeV )

    • CDet ( Peter ) : met with Brad and Bodgan , start testing all TDC modules in the Fastbus crates, figure out if we have all the cables, test cables we have on hand. Cover at least half of detector 1200 channels of TDCs.
    • GEM SSP readout (Danning / Ben ) : issue with MPD firmware fixed, need SSP firmware update : first version tested and had issue, EEL clean room : large setup ready to test (28 + 7 = 35 MPDs)

MPD current draw : need to have number to make sure crate can provide enough 5V current

    • GRINCH ( Bradley ) : getting started with CODA3, DAQ with cosmics no crash lately, might happen for longer run and higher rate, still want to update to latest VETROC firmware and driver
    • hodoscope ( Rachel ) : V1190 TDC setup , CODA 2.6.2 setup, add a V792 for calibration, setup detector to look at real signal
    • Bigbite ( Mark ) : need to get started on CODA3 and TS, need to make sure TS CPU boots up.
  • CAEN HV : try CAEN software, check with Steve to see what they have in Hall C
  • GRINCH trigger : with VTP/GTP at level of at least 200 ns might fit in 500 ns delay line

January 24th 2020

  • Updates
    • HCAL status : second half cable and working, some TDC not firing with FADC signal ( just for new half ), new scintillator trigger paddle, energy threshold study with cosmics, working on timewalk correction, need PMT from calibration soon. FADC could induce DC offset. will check with scope
    • GEM INFN : INFN DAQ still needs to check the isssue with MPD 17
    • GEM SSP readout ( Ben ) : issue with MPD firmware , Paolo had to fix, now working with new data format , being tested right for 12 MPDs, scaling to 32.
    • GRINCH DAQ crashing : no progress - need Bradley to contact Bryan
    • Bigbite ( Mark ) : VXS crate installed for TS
  • Gmn DAQ tasks : need to be ready January 2021
    • GEM readout : SSP ( VTP ?) about 1 month February
    • Switch CODA3 HRS, Bigbite, GRINCH, HCAL ( next two weeks )
    • TS for bigbite ( try finish end of January )
  • GEp DAQ tasks
    • geometrical correlation ECAL HCAL
    • HCAL trigger two crates
    • Fast clear GEM
  • Bigbite no Fastbus
    • No Edet, only shower on Fastbus : 243 channels = 16 FADC
    • VXS and VTP available from SoLID
    • if no FADC in HRS and no Compton : 3 FADC HRS - 2 FADC compton - currently have 15 total with FADC on loan from electronics group - SoLID needs 100 modules

December 13th 2019

  • GEp DAQ layout


  • GMn / Bigbite DAQ layout

  • Updates
    • INFN GEM

Isue with MPD 7 , when masked issue goes to another MPD

    • HCAL : all channels cabled, 1st half all issues fixed. Setting up for cosmics, test turn on all FADC and TDCs did not crash !
    • UVA GEM : old SSP firmware running with one SSP, ready testing new SSP firmware ( 16 MPDs ) , use cosmics

November 22th 2019

  • HCAL status
    • debugging old half, 6 channels have issues - TDC or discriminator issues maybe for 4 of them - 2 might be related to VME slot ? -

look with Bryan's decoder located in /site/coda/contrib/devel/dataDecode - if really looks like hardware issue can use VME bus analyzer

    • Ben setup the FADC gui - will be useful for FADC/detector debugging - Scott will write some documentation on wiki
    • start cabling other half of detector maybe next week
    • control pulser - use L1A to trigger LED, or use a VME pulser ( will try to give a DVCS LED board to Juan Carlos )
  • Grinch
    • DAQ issues - crashing for long runs
    • need to investigate - maybe check event size
  • BigBite
    • need new TS crate to check the final timing
    • takes about to generate Shower trigger : 60 ns cable + 60 ns logic + cable 25 m = 100 ns ~ 250 ns

October 25th 2019

  • Detectors DAQ status
    • BigBite
      • ADCs working fine running with CODA 2
      • signals from all shower preshower channels
      • can look at data with scripts
      • add TDCs to setup (27x7 + 2x27 ) = 243 channels, only sums so 52 channels ( 1 x 1877S )
      • try CODA3
    • HCal
      • FADC and F1 all running on HCAL
      • resolution checks, not good right now, might be due to discriminator ( Phillips 715 )
      • looking at cosmics
      • all cables for second half of HCAL
      • still unlock error on 1 TDC , need try lower resolution
      • still on CODA2
  • CODA3 testing (Bob) pdf
    • decoder in Podd can figure out if CODA3 or CODA2 data so should be transparent to SBS online library
    • trigger data not decoded yet
    • Eric tried CODA3 at beginning but still need some work for Fastbus
    • Need CODA3 for VTP
  • ATOM CPU testing
    • does not work out from the box

October 4th 2019

  • Bigbite DAQ status ( Eric )
    • slide here pdf
    • taking data but timing needs to be done to look at detector signals
    • Eric should look at existing PODD classes for shower to decode Fastbus
    • Shower trigger  : less than 200 ns
    • test trigger lines : 190 ns might

  • HCAL DAQ status ( Scott )
    • 2 crates : 1 VXS crates with FADC and 1 VME64X with F1 and FADC
- see cosmics in both crates in FADC , missing in TDC ( need to check new timing with 2 crates )
- resolution lock issue with F1, might be overheating
- add blanks to improve cooling
- reduce resolution of F1 to see if problem goes away
- latency from TI ( OT2 is prompt output less delay down to 200 ns)

  • UVA GEM status (Bryan / Kondo )
- 3 VME crates in clean room with TI ( 2 VME and 1 VXS ) : 2 planes per crate , 1 TI and one fiber, need 1 TI with two transceiver
- using CODA3
- wiki with information on wiki : SBS_DAQ:_GEM_EEL_Cleanroom_setup
- computer ready soon
- 3rd (VXS) crate with SSP and MPDs

  • SSP readout ( Ben )
- 16 MPD for 1 SSP to be tested
- all transceiver and fibers on site
- test radiation hardness of transceivers and fibers

  • BCM digital :
- daughter board design done, need to make prototypes and test

  • Certification of electronics : need to check with Ed Folts ( Nilanga question )

  • CDet
    • All TDC for coordinate in weldment : need checks ?
    • Software : running DAQ CODA and root scripts to analyze, need to check if other functionalities needed. Library can be loaded in Hall A / SBS analyzer
    • amount of TDCs for CDet sufficient
    • 25 TDCs - need to check with Bob